Hi i just have three problems for my homework. could you please help me?
1.) A group of people would like to buy a vacation cabin for $600,000 sharing the cost equally. If they could find 5 more people to join them, each person's share would be reduced by $6000. How many people are in the group?
2.) Write an equation in general form for the line passing through (-7,-10) and parallel to the line whose equation is 4x+2y-5=0
3.) A person standing on the roof of a building and throw a ball directly upward. The ball misses the rooftop on its way down and eventually strikes the ground. The function S(t)=-16t2+64t+80 descrobes the ball's height above the ground, S(t) in feet. T seconds after it was thrown

Answer by godeapurva(7)

Answer for your question no 1
Cost of the cabin is 600000$
Let the no of people be X , now they have mentioned that if 5 more people join the sharing each will save 6000$
So the amount contributed by each of X members is obviously > 6000$
And for that the number of contributors should also be less.
Now take 2 values 20 & 25
if x is 20 then contribution is 600000/20 = 30,000
and x+5 that is 25, then 600000/25 = 24,000
thus 30000-24000= 6000$
This fits our requirement.
Ans 25 people


Another method is
Let no of contributors = x
and contribution made = y
thus x*y = 600000
now (x+5)(y-6000)=600000
Now solve to get values for x and y
This will take much longer time than the 1st one.